Many states assess the skills of their students in various grades. One program that is available for this purpose is the National Assessment of Educational Progress (NAEP). One of the tests provided by the NAEP assesses the reading skills of twelfth-grade students. In a recent year, the national mean score was 285 and the standard deviation was 39 . Assume that these scores are approximately Normally distributed, N(285 , 39 ).

Use technology to find ( ±±0.0001) the proportion of students who have scores less than 318 .

.

Use technology to find ( ±±0.0001) the proportion of students who have scores greater than or equal to 318 .

.
Answer #1

Given ,

= 285 , = 39

We convert this to standard normal as

P(X < x) = P(Z < ( x - ) / )

a)

P(X < 318) = P(Z < (318 - 285) / 39)

= P(Z < 0.8462)

= 0.8013

b)

P(X >= 318) = 1 - P(X < 318)

= 1 - 0.8013

= 0.1987